# Using Causality for Enhanced Prediction of Web Traffic Time Series

### v1

@@ 4.2. Backbone Time Series Model
The Backbone Time Series Model, denoted as TS_Model, processes the input time series embedding $\mathbf{X}\in\mathbb{R}^{B\times N\times L_{x}\times C_{in}}$ and outputs a feature representation $h_{ts}^{iter}\in\mathbb{R}^{B\times N\times d_{ts}}$:
$h_{ts}^{iter}=\text{TS\_Model}(\mathbf{X}),$
which is concatenated with $h_{ccm}^{iter}$, generated by the CCMPlus in training or testing mode, to jointly predict the web service traffic time series.
Based on the baseline evaluation performances across three real-world web service traffic datasets at multiple prediction granularities (Tables 1, 2, 3, and 4) and insights from recent research [34; 46], TimesNet [48] and iTransformer [25] emerge as two strong-performing baselines among state-of-the-art time series models.
@@
TimesNet [48] identifies and utilizes multi-periodicity, decomposing temporal variations into intraperiod and interperiod components. The core module, TimesBlock, adaptively discovers periodicities and extracts features using a parameter-efficient inception block. iTransformer [25] repurposes the Transformer architecture for time series forecasting by applying attention and feed-forward networks on inverted dimensions. It embeds time points as variate tokens, allowing attention to capture multivariate correlations and the feed-forward network to learn nonlinear variate-specific representations.
@@
4.3. Optimization
For convenient combination with the Backbone Time Series Model and to improve generalization, the CCMPlus representation $h_{ccm}^{iter}$ is concatenated with the Time Series Model feature representation $h_{ts}^{iter}$. In general, the whole procedure can be formalized as follows:
$\hat{t}$ $=\text{MLP}\left(h_{ccm}^{iter}\mid h_{ts}^{iter}\right),$
(5) MSE Loss $=\frac{1}{B\cdot N\cdot L_{pred}}\sum_{b=1}^{B}\sum_{n=1}^{N}\sum_{l=1}^{L_{pred}}\left(t_{b,n,l}-\hat{t}_{b,n,l}\right)^{2},$
where $h_{ts}^{iter}\in\mathbb{R}^{B\times N\times d_{ts}}$, and both the ground truth $t$ and predicted values $\hat{t}\in\mathbb{R}^{B\times N\times L_{pred}}$. $B$ represents the batch size, $N$ denotes the number of web services within each batch, $d_{ts}$ refers to the representation dimension of the Backbone Time Series Models, and $L_{pred}$ specifies the prediction length of the web services in the time series.
5. Experiments
5.1. Datasets
We conduct experiments on three publicly available real-world web service traffic datasets from Alibaba Group  [^2] https://github.com/alibaba/clusterdata/blob/master/cluster-trace-microservices-v2022, Microsoft Azure  [^3] https://github.com/Azure/AzurePublicDataset/tree/master, and Ant Group  [^4] https://huggingface.co/datasets/kashif/App_Flow/viewer/. The Ant Group Traffic dataset includes 113 web services spanning a time range of 146 days. The Microsoft Azure Traffic dataset consists of 1,000 web services with a time range of 14 days. Similarly, the Alibaba Group Traffic dataset contains 1,000 web services, covering a total duration of 13 days. The datasets utilized in this study are available at: https://github.com/******.
5.2. Experiment Settings
5.2.1. Baselines
We evaluate our methods against the following baselines:
  • Large Language Models: Llama3 [42] and TimeLLM [20].
These baselines represent the SOTA approaches in their respective categories. Unlike these baselines, which ignore causal relationships among web services, our CCMPlus module explicitly incorporates them, enhancing feature representations and improving web service traffic prediction performance.
5.2.2. Evaluation Metrics
We evaluate model performances using Mean Squared Error (MSE) and Mean Absolute Error (MAE):
$\begin{array}[]{cc}\text{MSE}=\frac{1}{k}\sum_{i=1}^{k}\left(y_{i}-\hat{y}_{i}\right)^{2},&\text{MAE}=\frac{1}{k}\sum_{i=1}^{k}\left|y_{i}-\hat{y}_{i}\right|,\end{array}$
where $y_{i}$ and $\hat{y}_{i}$ denote the true and predicted values, respectively, and $k$ is the number of test samples in the test set.
5.2.3. Implementation Details
The implementation uses PyTorch 2.4.0 with random seeds {0, 2, 4}. Llama3 (8B) is fine-tuned using LoRA (rank 16, $\alpha=32$). The momentum value $m$ is 0.5. The batch size $B$ is 8, $C_{in}=16$, $\tau_{w}=100$, and the Adam optimizer is configured with a learning rate of 0.000001. The input length $L_{x}$ is 168, and the prediction length $L_{pred}$ is 1. Training is performed on an H100 GPU with $C_{out}=32$ and $\tau=[1,2,3,4]$. The model is trained for 15 epochs with an early stopping patience of 5 epochs based on validation performance.
5.2.4. Research Questions
The following research questions are investigated: 1) What are the overall performances and prediction accuracies of the models across different granularities? (Section 5.3) 2) How does the CCMPlus module impact prediction performance? (Section 5.4) 3) How does evaluation speed vary across methods? (Section 5.5) 4) How do the methods perform under varying hyperparameter settings? (Section 5.6)
5.3. Prediction Performance Analysis
Table 1. Average testing performance over three runs with a 30 minutes time interval across three datasets: Alibaba Group Traffic (124,000 test samples), Azure Traffic (134,000 test samples), and Ant Group Traffic (158,313 test samples).
@@ 30 Minutes Alibaba Group Traffic Microsoft Azure Traffic Ant Group Traffic Overall Mean
CCM+TimesNet (ours) 3.0206 ↓5.14% 0.5200 ↓3.81% 14.9237 ↓10.2% 0.6791 ↓5.4% 1.2897 ↓8.51% 0.9350 ↓6.40% 6.4113 0.7114
Table 2. Average testing performance over three runs with a 15 minutes time interval across three datasets: Alibaba Group Traffic (249,000 test samples), Azure Traffic (268,000 test samples), and Ant Group Traffic (316,739 test samples).

### v2

@@ 4.2 Backbone Time Series Model
The Backbone Time Series Model, denoted as BTSM, processes the input time series embedding $\mathbf{X}\in\mathbb{R}^{N\times L\times C}$ and outputs a feature representation $\mathbf{\widetilde{h}_{ts}}\in\mathbb{R}^{N\times Q}$: $\mathbf{\widetilde{h}_{ts}}=\text{BTSM}(\mathbf{X})$, which is concatenated with $\mathbf{\widetilde{h}_{ccm}}$ in the training or testing mode, to jointly predict the web service traffic time series.
@@
Based on the baseline evaluation performances across three real-world web service traffic datasets at multiple prediction granularities (Tables 1, 2 and Tables 5, 6 in Appendix) and insights from recent research [28, 41], TimesNet [43] and iTransformer [22] emerge as two strong-performing baselines among state-of-the-art time series models. TimesNet identifies and utilizes multi-periodicity, decomposing temporal variations into intraperiod and interperiod components. The core module, TimesBlock, adaptively discovers periodicities and extracts features using a parameter-efficient inception block. iTransformer repurposes the Transformer architecture for time series forecasting by applying attention and feed-forward networks on inverted dimensions. It embeds time points as variate tokens, allowing attention to capture multivariate correlations and the feed-forward network to learn nonlinear variate-specific representations.
@@
4.3 Optimization
For convenient combination with the Backbone Time Series Model and to improve generalization, the CCMPlus representation $\mathbf{\widetilde{h}_{ccm}}$ is concatenated with the time series model feature representation $\mathbf{\widetilde{h}_{ts}}$. In general, the whole procedure can be formalized as: $\hat{x}=\text{MLP}\left(\mathbf{\widetilde{h}_{ccm}}\mid\mathbf{\widetilde{h}_{ts}}\right)$, where $\hat{x}$ is the predicted time series value, and MLP denotes the multilayer perception that projects the concatenated hidden feature to the final prediction. Following previous work [22, 43], we employ the mean square error as the loss function to optimize the model parameters. The overall training and inference algorithm can be found in Appendix Section B.
Table 1: Prediction performances averaged over three runs with prediction granularity $\alpha$ set as 30 minutes. The best result is marked in bold. The $t$-test conducted on both metrics indicates that the improvement is statistical significant (p-value < 0.001).
@@ 30 Minutes Alibaba Group Traffic Microsoft Azure Traffic Ant Group Traffic Overall Mean
CCM+TimesNet (ours) 3.0206 ↓5.14% 0.5200 14.9237 0.6791 1.2897 ↓8.51% 0.9350 ↓6.40% 6.4113 0.7114
Table 2: Prediction performances averaged over three runs with prediction granularity $\alpha$ set as 5 minutes. The best result is marked in bold. The $t$-test conducted on both metrics indicates that the improvement is statistical significant (p-value < 0.001).
Method MSE MAE MSE MAE MSE MAE MSE MAE
Llama3 [38] 7.3436 1.1474 11.5244 1.3587 3.3047 1.5606 7.3909 1.3556
TimeLLM [16] 3.2588 0.5071 6.5393 0.5063 1.4974 1.0459 3.7652 0.6864
TimesNet [43] 2.1681 0.3925 2.8696 0.3527 1.3923 0.9822 2.1433 0.5758
CCM+TimesNet (ours) 1.8103 ↓16.50% 0.3394 ↓13.53% 2.6347 ↓8.19% 0.3150 ↓10.69% 1.2871 ↓7.56% 0.9287 ↓5.45% 1.9107 ↓10.85% 0.5277 ↓8.35%
5 Experiments
5.1 Datasets
We conduct experiments on three publicly available real-world web service traffic datasets from Ant Group [^4] https://huggingface.co/datasets/kashif/App_Flow/, Microsoft Azure [^5] https://github.com/Azure/AzurePublicDataset, and Alibaba Group [^6] https://github.com/alibaba/clusterdata. The Ant Group Traffic dataset includes 113 web services spanning a time range of 146 days. The Microsoft Azure Traffic dataset consists of 1,000 web services with a time range of 14 days. Similarly, the Alibaba Group Traffic dataset contains 1,000 web services, covering a total duration of 13 days.
5.2 Experiment Settings
5.2.1 Baselines
We evaluate our methods against the following baselines:
  • Large Language Models: Llama3 [38] and TimeLLM [16].
More details about baselines can be found in the related work (§2). These baselines represent the SOTA approaches in their respective categories. We apply the same experiment settings on these baselines and our methods to ensure fair comparison.
5.2.2 Evaluation Metrics
We evaluate the model performances using Mean Squared Error (MSE) and Mean Absolute Error (MAE):
$\begin{array}[]{cc}\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2},&\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|,\end{array}$
where $y_{i}$ and $\hat{y}_{i}$ denote the true and predicted values, respectively, and $n$ is the number of samples in the test set.
5.2.3 Implementation Details
The Llama3 (8B) baseline is fine-tuned using LoRA (rank is set as 16). The $\tau_{w}$ is set as 100, and $\tau\in[1,2,3,4]$. The momentum value $\lambda$ is 0.5. The channels of input and output of the convolution network are $C$ and $D$, $C$ and $D$ are set as $16$ and $32$, respectively. The batch size $B$ is 8, and the Adam optimizer is configured with a learning rate of 0.000001. The input length $L$ is 168. Training is performed on an H100 GPU. The model is trained for 15 epochs with an early stopping patience of 5 epochs based on validation performance. The prediction granularity parameter, $\alpha$, corresponds to the prediction horizon, distinguishing between long-term and short-term forecasts. $\alpha$ has following values: 1, 5, 15 and 30.
5.3 Prediction Performance Analysis
Performance Comparisons. To evaluate the performance of CCMPlus (CCM+), we compare two variants of our proposed framework, CCM+TimesNet and CCM+iTransformer, against the baselines. As shown in Table 1, we report the prediction results on three datasets, with the prediction granularity $\alpha$ (explained in Eq. 1) set to 30 minutes. Among all the baselines, Llama3 performs the worst. While large language models have demonstrated effectiveness in reasoning over sequential data, web traffic exhibits significantly higher volatility, posing substantial challenges for plain LLMs to accurately forecast future values. In contrast, TimeLLM inherently treats time series patches as tokens and employs task-specific prompt embeddings for forecasting. Although it captures patch correlations, it still underperforms on dynamic web traffic data compared to methods specifically tailored for time series analysis. Magicscaler and OptScaler, in particular, are designed for service workload prediction. While these models account for the volatility inherent in traffic time series, they fail to explicitly capture the underlying seasonal and trend components, which are crucial for predicting web service traffic driven by human behavior.
TimeMixer addresses this limitation by decomposing time series into seasonal and trend components across multiple periodicities, enabling it to learn decomposed temporal patterns ranging from fine-grained to macro-level perspectives. As a result, TimeMixer achieves superior prediction performance compared to both LLM-based methods and specialized workload prediction models. TimesNet leverages Fourier transforms to derive more adaptive periodicity terms, further reducing prediction errors. On the other hand, iTransformer treats independent time series as tokens, learning both temporal and cross-dimensional correlations by modeling token sequences, and achieves performance comparable to TimesNet. However, all the above methods focus solely on internal temporal patterns while neglecting external causal relationships across multiple time series. To address this limitation, we integrate CCMPlus with TimesNet and iTransformer—the two best-performing baseline models—to create two new variants. As observed, while TimesNet and iTransformer employ carefully designed architectures and achieve SOTA performance, integrating CCMPlus leads to further improvements. This highlights the importance of capturing the causality inherent in web service traffic to accurately predict traffic volume. Furthermore, CCMPlus effectively models this causality, contributing to enhanced predictive performance.
Table 3: Hyperparameter analysis of $C$, $D$, and $\tau$ on the Microsoft Azure Traffic dataset with 5-minute prediction granularity using CCM+TimesNet. The results are averaged over three experiments.
$C$ MSE MAE $D$ MSE MAE $\tau$ MSE MAE
12 2.6873 0.3178 28 2.8307 0.3149 [1, 2, 3] 2.7945 0.3157
14 2.8256 0.3385 30 2.6894 0.3223 [1, 2, 3, 4] (ours) 2.6347 0.3150
18 2.7078 0.3170 34 2.6357 0.3143
20 2.6311 0.3117 36 2.7461 0.3273
Prediction Granularity Analysis. Prediction granularity is an important factor for traffic forecasting. To assess this, we vary the prediction granularity $\alpha$ in the time series in {1, 5, 15, 30} minutes and compare the model performances in Table 6, 2, 5 and 1, respectively. (results for $\alpha=1$ and $\alpha=15$ are in Appendix §C.) Note that the more fine-grained time interval setting would generate more data samples. First, we observe that CCMPlus consistently improves the SOTA models (TimesNet and iTransformer), underscoring the importance of considering causality among service traffic patterns. Notably, CCMPlus demonstrates flexibility in integrating with various SOTA backbone models, highlighting its potential for performance enhancement. Second, the performance improvements diminish for the 1-minute prediction granularity setting. This is because service traffic is highly dynamic and volatile, and fine-grained time series data often lack a sufficient time duration to accumulate information necessary for capturing reliable causal relationships. Nevertheless, CCMPlus still achieves the best performance, further demonstrating its superiority in extracting hidden causal relationships and boosting prediction accuracy even under challenging conditions. We conduct a $t$-test [4] under both metrics shows that the improvement of our method over iTransformer and TimesNet is significant ($p$-value < 0.001).
5.4 Ablation Study
Table 4: Ablation study on the Microsoft Azure Traffic dataset with time granularity $\alpha$ set as 5 minutes.
Method MSE MAE
w/o BTSM 2.9981 0.3627
w/o MME&CER 2.8696 0.3527
w/o CER 2.7329 0.3320
CCM+TimesNet (ours) 2.6347 0.3150
The proposed model consists of the CCMPlus module (§4.1) and BTSM module (§4.2), where the CCMPlus encompasses of two key internal representations, i.e., the multi-manifold embedding and causality enhanced time-series representation (① and ④ introduced in §4.1). To assess the contribution of these modules and representations, we conduct a detailed ablation study using CCM+TimesNet on the most challenging dataset, Microsoft Azure Traffic. Table 4 presents the impact of model variants built on these modules and representations, yielding the following insights: (1) Excluding the Backbone Time Series Model (w/o BTSM) reduces performance. The deep learning-based backbone extracts essential seasonal and trend-related temporal features, which are crucial for accurate web service traffic forecasting. (2) By comparing the results when both Multi-Manifold Embedding (MME) and Causality Enhanced Representations (CER) are removed with those when only CER is removed, we can observe that MME provides diverse time-sensitive representations from multiple shadow manifold spaces formed by different time lags and embedding dimension values, which can lead to performance improvement for the advanced BTSM model; (3) Our complete model (i.e., BTSM+MME+CER) performs better than the variant without CER, which demonstrates that CER can generate feature representations that capture causal relationships among web services, enhancing traffic prediction accuracy.
5.5 Hyperparameter Tuning
Table 3 presents the hyperparameter analysis, evaluating the impact of the following key parameters: (1) $C$ defines the embedding dimension of the input time series $\hat{\mathbf{X}}$. As shown in Table 3, the model maintains stable performance across values, demonstrating robustness. Increasing this dimension enhances feature richness, but for efficiency, we set $C=16$. (2) $D$ defines the embedding dimension of shadow manifold points. A small value may miss features, while a large one can introduce noise. To balance representation quality and efficiency, we set $D=32$. (3) The values in $\tau$ define the dilation parameter in Conv1D, controlling the number of shadow manifolds. Increasing this number enhances feature extraction but incurs higher computational costs. To balance accuracy and efficiency, we set $\tau=[1,2,3,4]$.
(a) The causal relationship heatmap of web service A.
(b) Traffic time-series plots of web service A and service B.
(c) Traffic time-series plots of web service A and service C.
Figure 5: The heatmap of web service A’s causal relationships and the plots of the service A, B and C with the highest and lowest causal relationship, respectively.
5.6 Case Study
Using the Alibaba Group Traffic dataset, we compute causal relationships between web services with the CCMPlus module. To visualize, we compare Service A with twenty randomly sampled services and plot the causal correlation heatmap (Figure 5a), where Service A exhibits a strong causal correlation with the third service, referred to as Service B, and the weakest causal correlation with the seventh service, referred to as Service C. To further analyze this relationship, we plot the traffic time series of Services A, B and C over the same period (Figure 5b and 5c). Figure 5b indicate the causality correlation between service A and B, where a increase of A prompt the lagged increase of B (as marked in green circle), which exhibits a high causality score. From Figure 5c, we can observe that A and C show no discernible correlation, and the traffic of service C is steady and independent with A, which is also aligned with the heatmap results.
6 Conclusion
We propose the CCMPlus module, which captures causal relationships between web services to improve web service traffic prediction. CCMPlus first constructs multi-manifold for each web service traffic time series. It then estimates target time series from the multi-manifold space of other web services. By comparing these estimations with the ground truth, a causal correlation matrix is computed and used to generate the CCMPlus representation, quantifying inter-service causal dependencies. This representation is concatenated with the time series model’s output, bridging the gap in causal inference and enhancing prediction performance. We evaluate CCMPlus-integrated models on real-world web service traffic datasets, which demonstrates that CCMPlus consistently improves web service traffic prediction.
Appendix
We offer some technical details and reproducibility-related information as supplementary materials to help readers understand and reproduce our model.
Appendix A Cross Convergence Mapping
Convergent Cross Mapping (CCM) theory is a classical algorithm proposed to detect causal relationships between species. Inspired by this algorithm and preliminary experiments where web services show latent causality, we leverage CCM to enhance traffic prediction. To help readers better understand the CCM, we provide more details in this section.
The exposition of CCM is often illustrated using the context of the Lorenz system. As depicted in Figure 6, the trajectory of the Lorenz system forms a manifold $M$ in the state space. This manifold $M$ consists of a collection of points that represent all possible states of the Lorenz system over time, with these points connected to create a structured geometric space. The manifold, also referred to as the attractor, encompasses all trajectories and potential states $\underline{m}(t)$ of the system. Each state $\underline{m}(t)$ corresponds to a point in $M$, represented by the coordinate vector $\underline{m}(t)=[X(t),Y(t),Z(t)]$.
Figure 6: The CCM theory is explained using the canonical Lorenz system. This figure exemplifies the manifold $M$ and its corresponding shadow manifold $M_{x}$.
Figure 7: The point $\underline{y}(t)$ in the manifold $M_{y}$ corresponds to the contemporaneous point in time $\underline{x}(t)$ in the manifold $M_{x}$.
Figure 8: Convergent predictability as the time series length increases, assuming $Y$ has a causal effect on $X$.
    $\hat{Y}(t)\mid M_{X}=\sum_{i=1}^{E+1}w_{i}Y(t_{i}),$
    where $w_{i}$ represents the weight based on the distance between $\underline{x}(t)$ and its $i{\text{-th}}$ nearest neighbor in $M_{x}$, and $Y(t_{i})$ are the contemporaneous values of variable $Y(k)$. The weights $w_{i}$ are determined by:
    $w_{i}=\frac{u_{i}}{\sum_{j=1}^{E+1}u_{j}},$
    where
Appendix B Overall Training Algorithm
Algorithm 1 presents the training algorithm for our CCMPlus module corresponding to the procedure introduced in §4.1, which produce the causality enhanced representation $\mathbf{\widetilde{h}_{ccm}}$ and causal correlation matrix $\widetilde{\mathbf{M}}^{z}$, respectively.
Algorithm 1 Algorithmic Procedure for the CCMPlus Module
[…]
2:Output $\mathbf{\widetilde{h}_{ccm}}$ and $\widetilde{\mathbf{M}}^{z}$
3:Initialize $\mathbf{E}$ as an empty list.
4:for each $\tau_{i}$ in $\tau$ do
5:  $E_{i}\leftarrow\left\lfloor\dfrac{\tau_{w}}{\tau_{i}}\right\rfloor$
6:  if $E_{i}$ is even then
7:   $E_{i}\leftarrow E_{i}-1$
8:  end if
9:  Append $E_{i}$ to $\mathbf{E}$
10:end for
11:for $z\leftarrow 0,\ldots,Z$ do
12:  $\mathcal{H}=[]$, $\mathcal{C}=[]$
13:  for $i\leftarrow 1\ldots n$ do
14:   $\tau_{i}\leftarrow\tau[i-1]$,  $E_{i}\leftarrow\mathbf{E}[i-1]$
15:   if $(L-\tau_{i}\times(E_{i}-1))<0$ then
16:     continue
17:   end if
18:   Reverse the order of $\mathbf{X}$ along the $L$ dimension. Then reshape $\mathbf{X}$ to $\bar{\mathbf{X}}\in\mathbb{R}^{N\times C\times L}$
19:   $\mathbf{X}_{conv}=\text{Conv1D}\bigl{(}\bar{\mathbf{X}};\,\text{kernel\_size}=E_{i},\text{dilation}=\tau_{i},\bigr{)}$, $\mathbf{X}_{conv}\in\mathbb{R}^{N\times D\times\bar{L}}$
20:   if $\text{training\_flag}=\text{True}$ then
21:     Reshape $\mathbf{X}_{conv}$ to $\mathbf{X}_{ccm}\in\mathbb{R}^{N\times\bar{L}\times D}$, $\mathbf{Y}\leftarrow\hat{\mathbf{X}}[:,-\bar{L}:]$, $\mathbf{Y}$} is reversed along the $\bar{L}$ dimension
22:     Compute the distance matrix $\mathcal{D}\leftarrow\text{PairwiseDistances}(\mathbf{X}_{ccm}\mathbf{X}_{ccm})$
23:     Find the nearest neighbor’s indices of each web service $\mathcal{I}\leftarrow\text{argsort}(\mathcal{D},\text{dim}=-1)[:,:,1:(D+2)]$
24:     $d\leftarrow\text{Gather}(\mathcal{D},\mathcal{I})$
25:     $u_{z}\leftarrow\exp\!\Bigl{(}-\,\dfrac{d}{d[:,:,0{:}1]+\epsilon}\Bigr{)}$, $w_{z}\leftarrow\dfrac{u_{z}}{\sum(u_{z})+\epsilon}$
26:     $Y_{nearest}\leftarrow\text{Gather}(\mathbf{Y},\mathcal{I})$, $\hat{Y}\leftarrow\sum(w_{z}\times Y_{nearest})$
27:     $corr\leftarrow\dfrac{\mathrm{Cov}(\hat{Y},\,\mathbf{Y})}{\sigma(\hat{Y})\,\sigma(\mathbf{Y})+\epsilon}$, $\mathbf{M}\leftarrow corr^{\top}$
28:     if $z$ >0 then
29:      $\widehat{\mathbf{M}}^{z}(i)\leftarrow\lambda*\widetilde{\mathbf{M}}^{z-1}+(1-\lambda)*\mathbf{M}$
30:     else
31:      $\widehat{\mathbf{M}}^{z}(i)\leftarrow\mathbf{M}$
32:     end if
33:   else
34:     $\widehat{\mathbf{M}}^{z}(i)\leftarrow\widetilde{\mathbf{M}}^{z-1}$
35:   end if
36:   $\widehat{\mathbf{X}_{ccm}}\leftarrow\text{LinearProjection}(\text{Transpose}(\mathbf{X}_{ccm}))$
37:   $\mathbf{h}_{ccm}(i)=\widehat{\mathbf{M}}^{z}(i)\cdot\widehat{\mathbf{X}_{ccm}}$
38:   Append $h_{ccm}(i)$ to $\mathcal{H}$; Append $\widehat{\mathbf{M}}^{z}(i)$ to $\mathcal{C}$
39:  end for
40:  $\mathbf{\widetilde{h}_{ccm}}\leftarrow\text{Mean}(\text{Stack}(\mathcal{H}),\text{dim}=0)$
41:  $\widetilde{\mathbf{M}}^{z}\leftarrow\text{Mean}(\text{Stack}(\mathcal{C}),\text{dim}=0)$
42:end for
Appendix C Additional Performance Comparison Results
In this part, we provide additional results of the prediction granularity experiments ($\alpha=1$ and $\alpha=15$). It show that with our proposed CCMPlus module, the SOTA models can achieve further improvements on various prediction granularity setting across all the datasets. The $t$-test also demonstrates the improvements are statistical significant.
Table 5: Prediction performances averaged over three runs with prediction granularity $\alpha$ set as 15 minutes. The best result is marked in bold. The $t$-test conducted on both metrics indicates that the improvement is statistical significant (p-value < 0.001).